How do I Solve an equation in standard form for y-4=-(x-1)? It was easy at first but then I got confused?
Answer by jim_thompson5910(35256)

y-4=-(x-1)

y-4=-x+1

y=-x+1+4

y=-x+5

x+y=5

So the answer is x+y=5